# Mapping the Cloud-Driven Atmospheric Dynamics & Chemistry of an Isolated Exoplanet Analog with Harmonic Signatures

### v1

@@ 5.3 High-Altitude Heating Source
Mapping the variability in the wavelengths (3.05 to 3.5 m) bracketing the high-altitude, rotational-vibrational CH4 band (m, see Figure 6, Fifth Row) results in harmonics distinct from and largely anti-correlated with the low-altitude CH4 bands (see Figure 6, Fourth Row). For example, the CH4 absorption feature between 38.0 and 38.5 hr on the low-altitude CH4 map is replaced by a bright spot, which we tentatively interpret as an emission feature in the upper atmosphere, potentially driven by auroral heating from electron precipitation (e.g., Sandel et al., 1982; Melin et al., 2011; Müller-Wodarg et al., 2012; Pineda et al., 2024; McCarthy et al., 2025; Nasedkin et al., 2025). In this case, the CH4 abundance appears variable and is influenced and in-phase with underlying cloud modulation, but now we detect the CH4 in emission rather than absorption.
Distinct, high-altitude CH4 emission also aligns with previous findings for SIMP J0136 based on the same NIRSpec/PRISM dataset explored in this work. K-means clustering identified unique lightcurve morphology near the CH4 -band (McCarthy et al., 2025), and a 265 K temperature inversion was inferred at pressures (bar) corresponding to the same wavelengths (Nasedkin et al., 2025). The necessary power to create such a strong temperature inversion from electron precipitation-driven auroral heating (which is similar in strength to the 300 K heating found for WISE J1935, Faherty et al., 2024), was computed to be W by Nasedkin et al. (2025) while Jupiter-like auroral processes could only produce auroral power on the scale of W.
As discussed in Nasedkin et al. (2025), other upper atmospheric heating processes such as the breaking of gravity waves (Freytag et al., 2010) and Joule heating associated with atmospheric electric currents have been proposed for brown dwarfs based on observations of these mechanisms in outer solar system planets (Cowley et al., 2004; Ingersoll et al., 2021; Brown et al., 2022). Further studies will be necessary to break the degeneracies between these potential heating sources.
@@ 6 Summary
To understand how mineral cloud modulation, disequilibrium carbon chemistry, and upper atmospheric heating (potentially aurora-driven) each contribute to the well-established NIR variability of the isolated L/T transition, planetary-mass object SIMP J0136, we analyzed time-resolved spectroscopy gathered from JWST NIRISS and NIRSpec instruments, collected h apart, and originally presented in Akhmetshyn et al. (2025) and McCarthy et al. (2025). Performing Bayesian, Fourier fitting via Imber, a Python code developed in Plummer & Wang (2022, 2023); Plummer et al. (2024), we found the dominant harmonic modes (and associated parameters) for each 0.05m bin. Using these inferred harmonic modes and Sonora Bobcat (Marley et al., 2021) atmospheric models, we computed vertical variability maps demonstrating the flux variability at each pressure level and for specific molecular bands (i.e., H2O, CO, and CH4). We summarize our findings:
@@
SIMP J0136’s spectral variability evolves significantly over a period of h (see §4.1.1, 4.1.2 and Figures 3 and 4). NIRISS/SOSS spectra show the highest variability () in H2O absorption bands (1.75 to 2.05 m and 2.5 to 2.8 m) with muted variability at other wavelengths. However, NIRSpec/PRISM spectra have the highest variability at wavelengths m (corresponding to wavelengths expected to be affected by iron and forsterite cloud modulation and including the H2O absorption band) and also the longer-wavelength continuum (3.75 to 4.2 m) corresponding to deep pressure levels. We interpret this variation as dynamic modulation of the forsterite cloud-forming region.
@@
Harmonics with periodicities of h () are inferred preferentially at pressure levels corresponding to bar (the suspected iron and forsterite cloud modulation region, see §4.1.1, §4.1.2, §5.1 and Figures 3 and 4). As odd harmonics are indicative of North/South surface map asymmetry and equatorial latitudes are known to be cloudier in brown dwarfs, the observed harmonics suggest such an asymmetry is primarily in the cloud-forming (and likely near-equatorial) regions.
@@
We detect unique variability, independent of mean pressure-independent variation, in H2O (S/Nmax = 14.0), CO (S/Nmax = 13.0), and CH4 (S/Nmax = 14.9) molecular absorptions bands (see Figure 6). We find modulation in the wavelengths associated with the suspected forsterite cloud layer to be anti-correlated with wavelengths corresponding to H2O and CO and correlated with low-altitude CH4 and upper atmospheric temperature (see §4.2.2, §5.2 and Figure 6). We suggest that cloud formation drives the temperature and chemical structure in the overlying atmosphere, leading in part to oxygen sequestration and disequilibrium carbon chemistry in the upper atmosphere.
@@
We infer harmonic modes associated with the high-altitude, CH4 rotational-vibrational band distinct from those found in lower-altitude CH4 band wavelengths (see §4.2.2, §5.3 and Figure 6). Although the periodicities in both inferred vertical variability maps are comparable, they are anti-correlated, with the lower-altitude presumed CH4 absorption features transitioning to emission features. We interpret this emission as upper atmospheric heating (potentially of auroral origin) with a periodicity set by the CH4 abundance variability, which appears to be influenced by forsterite cloud modulation (as discussed in §5.2).
@@ Acknowledgements
M.K.P., F.P.C., and P.A.K. would like to thank the United States Air Force Academy Department of Physics and Meteorology, Martinson Honors Program, Space Physics & Atmospheric Research Center, and Center for Space Situational Awareness Research for supporting and enabling this research.

### v3

@@ 5.3 High-Altitude Heating Source
Mapping the deviations in the wavelengths (3.05 to 3.5 m) bracketing the high-altitude, rotational-vibrational CH4 band (m, see Figure 6, Fifth Row) results in harmonics distinct from and largely anti-correlated with the low-altitude CH4 bands (see Figure 6, Fourth Row). For example, the CH4 absorption feature between 38.0 and 38.5 hr on the low-altitude CH4 map is replaced by a bright spot, which we tentatively interpret as an emission feature in the upper atmosphere, potentially driven by auroral heating from electron precipitation (e.g., Sandel1982; Melin2011; Muller-Wodarg2012; Pineda2024; McCarthy2025; Nasedkin2025). In this case, the CH4 abundance appears variable and is influenced and in-phase with underlying cloud modulation, but now we detect the CH4 in emission rather than absorption.
Distinct, high-altitude CH4 emission also aligns with previous findings for SIMP J0136 based on the same NIRSpec/PRISM dataset explored in this work. K-means clustering identified unique lightcurve morphology near the CH4 -band (McCarthy2025), and a 265 K temperature inversion was inferred at pressures (bar) corresponding to the same wavelengths (Nasedkin2025). The necessary power to create such a strong temperature inversion from electron precipitation-driven auroral heating (which is similar in strength to the 300 K heating found for WISE J1935, Faherty2024), was computed to be W by Nasedkin2025 while Jupiter-like auroral processes could only produce auroral power on the scale of W.
As discussed in Nasedkin2025, other upper atmospheric heating processes such as the breaking of gravity waves (Freytag2010) and Joule heating associated with atmospheric electric currents have been proposed for brown dwarfs based on observations of these mechanisms in outer solar system planets (Cowley2004; Ingersoll2021; Brown2022). Further studies will be necessary to break the degeneracies between these potential heating sources.
@@ 6 Summary
To understand how mineral cloud modulation, disequilibrium carbon chemistry, and upper atmospheric heating (potentially aurora-driven) each contribute to the well-established NIR variability of the isolated L/T transition, planetary-mass object SIMP J0136, we analyzed time-resolved spectroscopy gathered from JWST NIRISS and NIRSpec instruments, collected h apart, and originally presented in Akhmetshyn2025 and McCarthy2025. Performing Bayesian, Fourier fitting via Imber, a Python code developed in Plummer&Wang2022; Plummer&Wang2023; Plummer2024, we found the dominant harmonic modes (and associated parameters) for each 0.05m bin. Using these inferred harmonic modes and Sonora Bobcat (Marley2021) atmospheric models, we computed vertical variability maps demonstrating the flux variability at each pressure level and deviations from bulk variability for molecular absorption features (i.e., H2O, CO, and CH4). We summarize our findings:
@@
SIMP J0136’s spectral variability appears to evolve over a period of h (see §4.1.1, 4.1.2 and Figures 3 and 4). NIRISS/SOSS spectra show the highest variability () in H2O absorption bands (1.75 to 2.05 m and 2.5 to 2.8 m) with muted variability at other wavelengths. However, NIRSpec/PRISM spectra have the highest variability at wavelengths m (corresponding to wavelengths expected to be affected by iron and forsterite cloud modulation and including the H2O absorption band) and also the longer-wavelength continuum (3.75 to 4.2 m) corresponding to deep pressure levels. We interpret this variation as dynamic modulation of the forsterite cloud-forming region.
@@
Harmonics with periodicities of h () are inferred preferentially at pressure levels corresponding to bar (the suspected iron and forsterite cloud modulation region, see §4.1.1, §4.1.2, §5.1 and Figures 3 and 4). As odd () harmonics are potentially indicative of North/South surface map asymmetry (depending on the limb darkening law) and equatorial latitudes are known to be cloudier in brown dwarfs, the observed harmonics may suggest such an asymmetry is primarily contained to the cloud-forming (and likely near-equatorial) regions.
@@
We detect deviations from the bulk (composite) variability for H2O (S/Nmax = 14.0), CO (S/Nmax = 13.0), and CH4 (S/Nmax = 14.9) molecular absorptions bands (see Figure 6). We find modulation in the wavelengths associated with the suspected forsterite cloud layer to be anti-correlated with wavelengths corresponding to H2O and CO and correlated with low-altitude CH4 and upper atmospheric temperature (see §4.2.2, §5.2 and Figure 6). We suggest that cloud formation drives the temperature and chemical structure in the overlying atmosphere, leading in part to oxygen sequestration and disequilibrium carbon chemistry in the upper atmosphere.
@@
We infer harmonic modes associated with the high-altitude, CH4 rotational-vibrational band distinct from those found in lower-altitude CH4 band wavelengths (see §4.2.2, §5.3 and Figure 6). Although the periodicities in both inferred vertical variability maps are comparable, they are anti-correlated, with the lower-altitude presumed CH4 absorption features transitioning to emission features. We interpret this emission as upper atmospheric heating (potentially of auroral origin) with a periodicity set by the CH4 abundance deviations, which appear to be influenced by forsterite cloud modulation (as discussed in §5.2).
@@ Acknowledgements
M.K.P., F.P.C., and P.A.K. would like to thank the United States Air Force Academy Department of Physics and Meteorology, Martinson Honors Program, Space Physics & Atmospheric Research Center, and Center for Space Situational Awareness Research for supporting and enabling this research. Furthermore, we would like to thank the anonymous reviewer for their thoughtful and detailed feedback and suggestions.
@@
This work is based on observations made with the NASA/ESA/CSA James Webb Space Telescope. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS for JWST. These observations are associated with programs GT 1209 (PI: Artigau) and GO 3548 (PI: Vos). The data described here may be obtained from the MAST archive at https://doi.org/10.17909/zryr-vs58 (catalog doi: 10.17909/zryr-vs58) and https://dx.doi.org/10.17909/T9RP4V (catalog doi:10.17909/T9RP4V).